OpenStudy (anonymous):

Topic: Math Find the value of the variable that satisfies the equation.Check solution. 1/3x+5=-4

OpenStudy (anonymous):

1/3x+5=-4 (multiplying 3 on both sides) (1/3x+5=-4)*3 x+15=-12 x=-12-15 x=-27

OpenStudy (skullpatrol):

Do you know how to check the solution?

OpenStudy (skullpatrol):

@jenny508 ?

OpenStudy (anonymous):

Yea